OpenStudy (anonymous):

Tan¤=5/12 find sin¤ [0;90]

OpenStudy (anonymous):

can you use your calculator to find the angle, using the inverse tan function ?

OpenStudy (anonymous):

This is a right triangle, so tan = opposite/adjacent = a/b and third side is hypotenuse, c^2=a^2 + b^2 with sin = opp/hyp = a/c your answer after calculation

OpenStudy (anonymous):

hey @douglaswinslowcooper , how did you know that the triangle is Right angled ? :)

OpenStudy (anonymous):

Because tan, sin, cos etc. are only defined for right triangles.

OpenStudy (anonymous):

hey @douglaswinslowcooper , in the link below, the triangle doesn't have to be right-angled, so u can apply Sin/Cos/Tan to any angled triangle http://www.mathsisfun.com/algebra/trig-sine-law.html May be here it depends what level is @Sarel - if they are working with right- angled triangles only then your answer is Spot-ON :)